what is the coordinate of the point of intersection for the two lines below? x-2y=-2, y=-6x+40
Answer by Mathtut(3670)

x-2y=-2 ......eq 1
y=-6x+40.......eq 2
:
take y's value from eq 2 and plug it into eq 1
:
x-2(-6x+40)=-2
:
x+12x-80=-2
:
13x=78
:
x=6
:
y=-6(6)+40=4
:
coordinate solution is (6,4)